Ask your own question, for FREE!
Mathematics 8 Online
OpenStudy (anonymous):

P(A) = 0.42. What are the odds for A?

OpenStudy (anonymous):

favourable cases / unfavourable cases ie. 0.42/1-0.42

OpenStudy (anonymous):

-0.16 ?

OpenStudy (anonymous):

negative sign?

OpenStudy (anonymous):

is 0.42/1-0.42 the equation?

OpenStudy (anonymous):

yeah but 1-0.42 = 0.58 and 0.42 = +0.42 so i can figure out wheres this negative sign come from..

OpenStudy (anonymous):

That was the wrong way to solve, but thanks anyways!!

OpenStudy (anonymous):

it was actually 0.42:0.58 = 42:58 =21:29

OpenStudy (anonymous):

hmm..thats what i told u.. (:

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!